OpenStudy (jordanloveangel):

Which answer shows this equation in standard form? 7 – 3(x – y) = –5x + 2 A. –8x + 3y = –5 B. –2x – 3y = 5 C. 2x + 3y = –5 D. 2x – 3y = –5

OpenStudy (anonymous):

\[ 7 - 3(x-y) = 7 - 3x + 3y = -5x + 2 \iff 5 + 2x + 3y = 0 \implies 2x + 3y = -5 \\ \boxed{\textbf{(C)}} \]

OpenStudy (jordanloveangel):

thanx @Ahaanomegas

OpenStudy (anonymous):

You're welcome!